# Lessons learned from field demonstrations of model predictive control and reinforcement learning for residential and commercial HVAC: A review

### v1

@@ 4.4 Experiment protocol choices
Fig. 6 shows the numbers of residential (darker bars) and commercial (lighter) field tests corresponding to each choice in the three experiment protocols. The left plot shows that the majority of both residential and commercial tests benchmarked against measurements. However, seven (13%) of 52 residential tests and 47 (33%) of 153 (one did not report) commercial tests benchmarked against simulations. The center plot shows that 38 (73%) of 52 residential tests controlled a whole building, whereas 110 (71%) of 154 commercial tests controlled a subset of building spaces. The right-most plot in Fig. 6 shows that 15 (29%) of 52 residential tests and 67 (46%) of 146 (eight did not report) commercial tests lasted less than seven days. Only 20 (38%) of 52 residential tests and 39 (27%) of 146 commercial tests lasted one month or more.
@@ 5.1 Benchmarking method
The choice of benchmark (also known as the baseline or counterfactual) influences the estimates of energy and cost savings from advanced HVAC control relative to a baseline control method. Ideally, the advanced control method and the baseline control method would be implemented in the same building under the same weather and occupancy conditions. However, this is only possible if researchers have access to two identical, side-by-side buildings, or to a building inside an environmental chamber where weather conditions can be repeatably emulated. Such experimental testbeds are nearly nonexistent, so researchers almost always must resort to less rigorous benchmarking methods, of which there are two broad categories. First, measurement-based benchmarks measure the field performance of the advanced control method and the baseline control method in the same building at different times, and therefore under potentially different weather and occupancy conditions. This approach typically involves adjusting for weather variations, for example, through normalization of energy use by heating degree-days or cooling degree-days. Second, simulation-based benchmarks compare the measured performance of the advanced control method in the real building to the simulated performance of the baseline control method in a computer model of the building, typically under the same weather conditions observed in the advanced control test period. Common model structures include thermal networks, time-series models, neural networks, regression trees, support vector machines, and high-fidelity building energy simulation programs such as EnergyPlus [83] and TRNSYS [84].
@@ 5.4 Energy and cost savings in filtered data
Applying the three experiment protocol filters described in this section – selecting for studies that used measurement-based benchmarking, controlled a whole building or a large subset of a building, and tested for at least seven days – left 25 (48%) and 28 (18%) tests from the initial residential and commercial pools of 52 and 154 field tests, respectively. These tests are attributed to 9 out of 24 (37%) and 21 out of 80 (26%) unique residential and commercial papers. We also excluded 16 of the 206 total tests due to incomplete reporting of experiment protocols. Fig. 10 shows the energy and cost savings over these remaining 53 tests. The energy savings (top row of plots) ranged from -0.4% to 60.4%. The duration-weighted average energy savings were 22% for residential buildings (left column) and 15% for commercial buildings (right). The cost savings (bottom) ranged from 1.3% to 39.0%. The duration-weighted average cost savings were 16% for residential buildings and 13% for commercial buildings.
@@ 6 Discussion
Here we discuss a number of potentially informative analyses we conducted that ended up being inconclusive due to small sample sizes or inconsistent reporting of relevant information (Section 6.1). Section 6.2 discusses what information we found on costs, labor, and other challenges related to deploying, operating, and maintaining MPC and RL systems. Section 6.3 recommends experiment protocols and research directions for future field work. Section 6.4 discusses limitations of this paper and opportunities for future literature reviews to address those limitations.
@@ 6.1 Inconclusive analyses
This section discusses our attempts to analyze how energy and cost savings from MPC and RL depend on the electricity rate structure, the building properties, the control architecture, and the climate zone and operational modes. Although studying these factors could help identify settings where MPC and RL are particularly effective, inconsistencies in reporting across studies made it difficult to draw robust conclusions.
@@ 6.1.2 Building properties
We hypothesize that building properties can substantially affect cost savings from MPC and RL. Thermal mass, for example, influences a building’s capacity to store thermal energy, and therefore a control system’s ability to shift energy use across time. Similarly, insulation level has a significant impact on HVAC performance by determining the extent of heat retention and thermal losses, directly impacting load-shifting potential. We attempted to test this hypothesis by categorizing buildings as low-to-moderate vs. moderate-to-high insulation levels and thermal masses (see Sections 4.3.4 and 4.3.5). However, few papers reported this information, making accurate classification of studies very challenging. After applying the experimental filters outlined in Section 5, sample sizes were too small to draw statistically significant conclusions.
@@ 6.1.3 Control architecture
We investigated whether the control architecture – for example, adjusting high-level setpoints vs. low-level actuator positions – influences energy and cost savings from MPC and RL (see Section 4.3.6). We attempted to test hypothesis that high- and low-level control architectures lead to different savings distributions. However, after applying the experiment protocol filters in Section 5, the sample sizes were too small to draw statistically significant conclusions.
@@ 6.1.4 Climate and operating mode
We investigated the influence of climate conditions and HVAC operating modes (heating, cooling, or both) on energy and cost savings from advanced control. Specifically, we categorize studies based on climate zones, heating and cooling degree-days, diurnal temperature variations, and operating modes (see Sections 4.2 and 4.3.3). However, after applying the experiment protocol filters in Section 5, no clear trends emerged, possibly due to inconsistencies in how studies reported climate data and operational modes, to variations in building properties and HVAC equipment across studies, or to the uneven geographic distribution of field studies.
@@ 6.2 Costs, labor, and other challenges
After applying the experiment protocol filters outlined in Section 5, the duration-weighted average cost savings were 16% in residential buildings and 13% in commercial buildings. While these cost savings are potentially attractive, making the economic case for advanced HVAC control also requires understanding the costs of deployment, operation, and maintenance. Only 13 papers provide information on these costs, which we organize into four categories: instrumentation, data and communication, labor, and cost-benefit analysis. This section collects what little cost information has been reported.
@@ 6.2.1 Instrumentation
Instrumentation costs cover expenses associated with devices for sensing, actuation, communication, and computing. In residential applications, Finck et al. [91] reported hardware costs including a data logger (\$2,800, about \$3,450 in 2025), a flow meter (\$200+, about \$246+ in 2025), and a smart meter (\$200+, about \$246+ in 2025). Wang et al. [92] described a low-cost plug-and-play system priced at \$50 (about \$52.5 in 2025) that used two smart wall sockets and two temperature sensors. Pergantis et al. [93] detailed an approach that employed a power meter (\$100) and smart thermostats with Wi-Fi control via manufacturer application programming interfaces, thereby avoiding custom hardware. For commercial applications, Goyal et al. [94] reported a \$215 (about \$280 in 2025) occupancy sensor for an occupancy-based setback controller, while Sturzenegger et al. [14] documented additional instrumentation costs – including new room temperature sensors, solar radiation sensors, and blinds actuators – totaling 3,000 to 6,000 CHF (about \$3,400 to \$6,800 in 2025). Ham et al. [95] noted that their MPC infrastructure cost only \$975 (about \$1,200 in 2025) in hardware. Ham et al. [95] also spent \$7,330 (about \$9,020 in 2025) on instrumentation for measurement and verification, a cost incurred primarily for research purposes rather than routine operation.
@@ 6.2.2 Data and networking
Data and networking costs include expenses for off-site data sources (such as weather forecasts) and on-site information technology services necessary to support reliable data access and communication. In residential settings, Pergantis et al. [93] used manufacturer-provided application programming interfaces, which cost nothing for their research purposes but might entail subscription fees for deployment at scale. In commercial settings, Sturzenegger et al. [14] reported an annual subscription to weather forecast data ranging from 100 to 600 CHF (about \$115 to \$680 in 2025).
@@ 6.2.3 Labor
Labor costs encompass the human effort required to deploy, operate, and maintain an advanced HVAC control system. In residential contexts, Pergantis et al. [16] reported that MPC deployment required 190 engineer-days in total, delineated between 150 days for one-time tasks (which would not have to be repeated for each new deployment) and 40 days for recurring tasks (which would have to be repeated for each new deployment) such as device installation, calibration, and troubleshooting. Ongoing maintenance activities are also labor intensive, as noted by Knudsen et al. [96], who stressed the need for periodic model recalibration, and Svetozarevic et al. [97], who emphasized the need for continuous commissioning. In commercial settings, Sturzenegger et al. [14] highlighted high labor demands for model development and for integrating and commissioning devices. Ruano et al. [98] described a one-week installation process for hardware setup and system integration, while Granderson et al. [99] reported that installation, information technology approvals, troubleshooting, and stability monitoring can require up to three days of staff time. Blum et al. [15] quantified these efforts in a commercial MPC deployment. They estimated 239 person-days total, with 70 person-days for preparation, 79 for model development and integration, 30 for controller software development, and 60 for deployment (commissioning, building management system integration, and maintenance).
@@ 6.2.4 Cost-benefit analysis
Project finance metrics – such as simple payback, benefit-to-cost ratio, net present value, internal rate of return, etc. – provide key information to businesspeople who decide whether to deploy advanced HVAC control systems. In residential buildings, Thorsteinsson et al. [100] estimated a cost recovery of €1,200 (about \$1,320 in 2025) over 10 years. For commercial applications, Ham et al. [95] reported a payback period of 13 heating months for MPC infrastructure, while Sturzenegger et al. [14] cautioned that high deployment and engineering costs may outweigh potential savings unless improvements in automation and modeling frameworks are achieved.
@@ 6.3.3 Report the context
    Whether the building is old or new; is insulated poorly or well; is drafty or tight; and is thermally light or heavy. Ideally, researchers would provide quantitative metrics for these building characteristics, such as the date of construction or renovation; the estimated thermal resistance of the building envelope; the estimated air changes per hour of outdoor air infiltration; and the estimated thermal capacitance of the building materials. If these data are unavailable, however, we recommend that researchers provide at least their own qualitative assessments of the building characteristics.
@@
    Characterization of the climate zone (e.g., hot, moderate, or cold; dry or humid).
@@ 6.3.4 Expand the hardware scope
Almost all of the reviewed field studies controlled HVAC systems only, ignoring interactions with other equipment. We recommend that future field demonstrations coordinate HVAC systems alongside other distributed energy resources, such as electric vehicles, stationary batteries, thermal storage, or solar photovoltaics. We conjecture that jointly coordinating a variety of distributed energy resources under one control system, rather than separately controlling each class of equipment, will substantially improve the overall business case for advanced control.
@@ 6.3.5 Explore emerging objectives
While many simulation studies have explored the potential of HVAC equipment to actively participate in power grid operations, few field studies have evaluated these opportunities in real, grid-interactive buildings. We recommend that future field studies control HVAC equipment (possibly alongside other distributed energy resources) to balance a range of objectives, including both traditional goals such as improving energy efficiency and emerging goals such as peak shaving, arbitraging day-ahead or real-time wholesale energy prices, and providing power grid reliability services such as spinning reserve, frequency regulation, and voltage support.
@@ 6.4 Limitations and future literature reviews
While this paper reviews all known peer-reviewed field demonstrations of MPC and RL for residential and commercial HVAC, it has several limitations. To keep the analysis tractable, we limited the scope of our quantitative analyses to peer-reviewed journal articles and conference papers, excluding white papers, theses, dissertations, and technical reports. Future literature reviews could augment our open-access dataset [19] to include these sources.
This paper also focused primarily on MPC and RL, as these are two popular algorithms for advanced HVAC control. Future work could expand the scope to include field demonstrations of other advanced control methods, comparing their benefits and costs to those of MPC and RL. Some of these methods include open-loop optimal control [101, 102, 103, 104], fuzzy logic control [105, 106], and well-designed rule-based control [107, 108, 109, 110, 111, 112], as well as robust control, adaptive control, and hybrid approaches that combine optimization and learning-based methods.
The analysis in this paper was also limited by sparse and inconsistent reporting on electricity rate structures, building properties, control architectures, climate data, and other contextual information. If future field studies standardize reporting of these data, future literature reviews could revisit the hypotheses discussed in Section 6.1 to identify contexts in which advanced HVAC control is particularly effective.
@@ 7 Conclusion
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this study, the authors used ChatGPT in order to improve the readability and language and to check spelling and grammar. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
@@
Arash J. Khabbazi: Conceptualization, Methodology, Software, Validation, Formal Analysis, Investigation, Data Curation, Writing - Original Draft, Writing - Review & Editing, Visualization, Supervision. Elias N. Pergantis: Methodology, Formal Analysis, Investigation, Writing - Original Draft, Writing - Review & Editing. Levi D. Reyes Premer: Formal Analysis, Investigation, Writing - Review & Editing. Panagiotis Papageorgiou: Formal Analysis, Investigation, Writing - Review & Editing. Alex H. Lee: Formal Analysis, Investigation, Writing - Review & Editing. James E. Braun: Conceptualization, Methodology, Writing - Review & Editing. Gregor P. Henze: Conceptualization, Methodology, Writing - Review & Editing. Kevin J. Kircher: Conceptualization, Methodology, Validation, Writing - Review & Editing, Supervision, Project Administration, Funding Acquisition.
@@
Data have been made publicly available in this repository [19].

### v6

@@ 4.5 Experiment protocol choices
Figure 6: Numbers of residential and commercial tests within each category of experiment protocol choices.
Fig. 6 shows the numbers of residential (darker bars) and commercial (lighter) field tests corresponding to each choice in the three experiment protocols. The left plot shows that the majority of both residential and commercial tests benchmarked against measurements. However, seven (14%) of 51 residential tests (one not benchmarked) and 47 (33%) of 153 (one not benchmarked) commercial tests benchmarked against simulations. The center plot shows that 38 (73%) of 52 residential tests controlled a whole building, whereas 110 (71%) of 154 commercial tests controlled a subset of building spaces. The right-most plot in Fig. 6 shows that 15 (29%) of 52 residential tests and 67 (46%) of 146 (eight did not report) commercial tests lasted less than seven days. Only 20 (38%) of 52 residential tests and 39 (27%) of 146 commercial tests lasted one month or more.
@@ 5.1 Benchmarking method
The choice of benchmark (also known as the baseline or counterfactual) influences the estimates of energy and cost savings from advanced HVAC control relative to a baseline control method. Ideally, the advanced control method and the baseline control method would be implemented in the same building under the same weather and occupancy conditions. However, this is only possible if researchers have access to two identical, side-by-side buildings, or to a building inside an environmental chamber where weather conditions can be repeatably emulated. Such experimental testbeds are nearly nonexistent, so researchers almost always must resort to less rigorous benchmarking methods, of which there are two broad categories. First, measurement-based benchmarks measure the field performance of the advanced control method and the baseline control method in the same building at different times, and therefore under potentially different weather and occupancy conditions. This approach typically involves adjusting for weather variations, for example, through normalization of energy use by heating degree-days or cooling degree-days. Second, simulation-based benchmarks compare the measured performance of the advanced control method in the real building to the simulated performance of the baseline control method in a computer model of the building, typically under the same weather conditions observed in the advanced control test period. Common model structures include thermal networks, time-series models, ANNs, regression trees, support vector machines, and high-fidelity building energy simulation programs such as EnergyPlus and TRNSYS.
@@
Both simulation- and measurement-based benchmarking methods introduce uncertainty into energy and cost savings estimates. Reporting confidence intervals alongside point estimates can characterize the degree of uncertainty present, and thereby inform readers about the trustworthiness of savings estimates. However, confidence intervals are rarely reported in field experiments. Among residential papers, only three out of 23 (one not applicable) included confidence intervals, while just six out of 77 (three not applicable) commercial papers did the same. Overall, only 9% of papers reported confidence intervals for their energy or cost savings estimates.
@@ 5.4 Energy and cost savings in filtered data
Applying the three experiment protocol filters described in this section – selecting for studies that used measurement-based benchmarking, controlled a whole building or a large subset of a building, and tested for at least seven days – left 25 (48%) and 28 (18%) tests from the initial residential and commercial pools of 52 and 154 field tests, respectively. These tests are attributed to 9 out of 24 (37%) residential papers and 21 out of 80 (26%) commercial papers, or 29% of the 104 total papers. We also excluded 16 of the 206 total tests due to incomplete reporting of experiment protocols. Fig. 10 shows the energy and cost savings over these remaining 53 tests. The energy savings (top row of plots) ranged from -0.4% to 60.4%. The duration-weighted average energy savings were 22% for residential buildings (left column) and 15% for commercial buildings (right). The cost savings (bottom) ranged from 1.3% to 39.0%. The duration-weighted average cost savings were 16% for residential buildings and 13% for commercial buildings.
@@ 6 Discussion
Here we discuss a number of potentially informative analyses we conducted that ended up being inconclusive due to small sample sizes or inconsistent reporting of relevant information (Section 6.1). We then summarize any available information on costs, labor, and other challenges related to deploying, operating, and maintaining MPC and RL systems (Section 6.2). Based on the reviewed studies, we recommend experiment protocols and research directions for future field work (Section 6.3). Lastly, we discuss limitations of this paper and opportunities for future literature reviews to address those limitations (Section 6.4).
@@ 6.1 Inconclusive analyses
In this section, we discuss our attempts to analyze how energy and cost savings from MPC and RL depend on the electricity rate structure, the building properties, the control architecture, and the climate zone and operational modes. Although studying these factors could help identify settings where MPC and RL are particularly effective, inconsistencies in reporting across studies made it difficult to draw robust conclusions.
@@ 6.1.2 Building properties
We hypothesize that building properties can substantially affect cost savings from MPC and RL. Thermal mass, for example, influences a building’s capacity to store thermal energy, and as a result, a control system’s ability to shift energy use across time. Similarly, insulation has a significant impact on HVAC performance by determining the extent of heat retention and thermal losses, which affect load-shifting potential. We attempted to test this hypothesis by categorizing buildings as low-to-moderate vs. moderate-to-high insulation levels and thermal masses (see Sections 4.3.4 and 4.3.5). However, few papers reported this information, making accurate classification of studies very challenging. After applying the experimental filters outlined in Section 5, sample sizes were too small to draw statistically significant conclusions.
@@ 6.1.3 Control architecture
We investigated whether the control architecture – for example, adjusting high-level setpoints vs. low-level actuator positions – influences energy and cost savings from MPC and RL (see Section 4.3.6). We attempted to test hypothesis that high- and low-level control architectures lead to different savings distributions. However, after applying the experiment protocol filters defined in Section 5, the sample sizes were too small to draw statistically significant conclusions.
@@ 6.1.4 Climate and operating mode
We investigated the influence of climate conditions and HVAC operating modes (heating, cooling, or both) on energy and cost savings from advanced control. Specifically, we categorized studies based on climate zones, heating and cooling degree-days, diurnal temperature variations, and operating modes (see Sections 4.2 and 4.3.3). However, after applying the experiment protocol filters defined in Section 5, no clear trends emerged. This is possibly due to inconsistencies in how studies reported climate data and operational modes, to variations in building properties and HVAC equipment across studies, or to the uneven geographic distribution of field studies.
@@ 6.2 Costs, labor, and other challenges
After we applied the experiment protocol filters defined in Section 5, the duration-weighted average cost savings were 16% in residential buildings and 13% in commercial buildings. While these cost savings are potentially attractive, making the economic case for advanced HVAC control also requires understanding the costs of deployment, operation, and maintenance. Only 13 papers provide information on these costs, which we organize into four categories: instrumentation, data and communication, labor, and cost-benefit analysis. This section collects what little cost information has been reported.
@@ 6.2.1 Instrumentation
Instrumentation costs include expenses associated with devices for sensing, actuation, communication, and computing. With a focus on residential buildings, Finck et al. [91] reported hardware costs including a data logger (\$2,800, about \$3,450 in 2025), a flow meter (\$200+, about \$246+ in 2025), and a smart power meter (\$200+, about \$246+ in 2025). Wang et al. [92] described a low-cost plug-and-play system priced at \$50 (about \$52.5 in 2025) that used two smart wall sockets and two temperature sensors. Pergantis et al. [93] detailed an approach that used a power meter (\$100) and smart thermostats with Wi-Fi control via manufacturer application programming interfaces. With a focus on commercial buildings, Goyal et al. [94] reported a \$215 (about \$280 in 2025) occupancy sensor for an occupancy-based setback controller. Sturzenegger et al. [16] documented additional instrumentation costs, such as new room temperature sensors, solar radiation sensors, and blinds actuators, costing 3,000 to 6,000 CHF (about \$3,400 to \$6,800 in 2025). Ham et al. [95] noted that their MPC infrastructure cost only \$975 (about \$1,200 in 2025) in hardware. They additionally spent \$7,330 (about \$9,020 in 2025) on instrumentation for measurement and verification, a cost mostly for research purposes rather than routine operation.
@@ 6.2.2 Data and networking
Data and networking costs include expenses for off-site data sources, such as weather forecasts, and on-site information technology services required for reliable data access and communication. Pergantis et al. [93] used manufacturer-provided application programming interfaces, which cost nothing for their research purposes but might entail subscription fees for deployment at scale. Sturzenegger et al. [16] reported an annual subscription to weather forecast data that ranged from 100 to 600 CHF (about \$115 to \$680 in 2025).
@@ 6.2.3 Labor
Labor costs include the human effort required to deploy, operate, and maintain an advanced HVAC control system. In residential buildings, Pergantis et al. [18] reported that MPC deployment required 190 engineer-days in total, delineated between 150 days for one-time tasks (which would not have to be repeated for each new deployment) and 40 days for recurring tasks (which would have to be repeated for each new deployment) such as device installation, calibration, and troubleshooting. Ongoing maintenance activities are also labor intensive, as noted by Knudsen et al. [96], who stressed the need for periodic model recalibration, and Svetozarevic et al. [97], who emphasized the need for continuous commissioning. In commercial buildings, Sturzenegger et al. [16] highlighted high labor demands for model development and for integrating and commissioning devices. Ruano et al. [98] described a one-week installation process for hardware setup and system integration. Granderson et al. [99] reported that installation, information technology approvals, troubleshooting, and stability monitoring can require up to three days of staff time. Blum et al. [17] quantified these efforts in a commercial MPC deployment. They estimated 239 person-days total, with 70 person-days for preparation, 79 for model development and integration, 30 for controller software development, and 60 for deployment (commissioning, building management system integration, and maintenance).
@@ 6.2.4 Cost-benefit analysis
Project finance metrics – such as simple payback, benefit-to-cost ratio, net present value, internal rate of return, etc. – provide key information to business managers who decide whether to deploy advanced HVAC control systems. In residential buildings, Thorsteinsson et al. [100] estimated a cost recovery of €1,200 (about \$1,320 in 2025) over 10 years. In commercial buildings, Ham et al. [95] reported a payback period of 13 heating months for MPC infrastructure. Sturzenegger et al. [16] cautioned that high deployment and engineering costs may outweigh potential savings unless improvements in automation and modeling frameworks are achieved.
@@ 6.3.3 Report the context
    Estimates of the date of construction or latest renovation, the thermal resistance of the building envelope, the outdoor air infiltration rate, and the thermal capacitance of the building materials. If these data are unavailable, we recommend that researchers provide qualitative assessments of whether the building is old or new; is insulated poorly or well; is drafty or tight; and is thermally light or heavy.
@@
    Characterization of the climate zone (e.g., hot, mixed, or cold; dry, humid, or marine).
@@ 6.3.4 Expand the hardware scope
Almost all of the reviewed field studies controlled HVAC systems only, ignoring interactions with other equipment. We recommend that future field demonstrations coordinate HVAC systems alongside other distributed energy resources (DERs), such as electric vehicles (EVs), stationary batteries, thermal storage, or solar photovoltaics (PV). We conjecture that jointly coordinating a variety of DERs under one control system, rather than separately controlling each class of equipment, will substantially improve the overall business case for advanced control.
@@ 6.3.5 Explore emerging objectives
While many simulation studies have explored the potential of HVAC equipment to actively participate in power grid operations, few field studies have evaluated these opportunities in real, grid-interactive buildings. We recommend that future field studies control HVAC equipment (possibly alongside other DERs) to balance a range of objectives, including both traditional goals such as improving energy efficiency and emerging goals such as peak shaving, arbitraging day-ahead or real-time wholesale energy prices, and providing power grid reliability services such as spinning reserve, frequency regulation, and voltage support.
@@ 6.4 Limitations and future literature reviews
While this paper reviewed all peer-reviewed field demonstrations of MPC and RL for residential and commercial HVAC of which the authors are aware, it has several limitations. To keep the analysis tractable, we limited the scope of our quantitative analyses to peer-reviewed journal articles and refereed conference papers, excluding white papers, theses, dissertations, and technical reports. Future literature review papers could expand our open-access dataset [20] to include these sources.
This paper also focused primarily on MPC and RL, two popular algorithms for advanced HVAC control. Future work could expand the scope to include field demonstrations of other advanced control methods, comparing their benefits and costs to those of MPC and RL. Some of these methods include open-loop optimal control [101, 102, 103, 104], fuzzy logic control [105, 106], rule-based control [107, 108, 109, 110, 111, 112], robust control, adaptive control, and hybrid approaches that combine optimization and learning-based methods. Furthermore, given the growing interest in developing model-free control methods, variants of MPC such as DeePC [113, 114, 115] could be investigated further.
The analysis in this paper was also limited by sparse and inconsistent reporting on electricity rate structures, building properties, control architectures, climate data, and other contextual information. If future field studies standardize reporting of these data, future literature review papers could revisit the hypotheses discussed in Section 6.1 to identify contexts in which advanced HVAC control is particularly effective.
@@ 7 Conclusion
Arash J. Khabbazi: Conceptualization, Methodology, Software, Validation, Formal Analysis, Investigation, Data Curation, Writing - Original Draft, Writing - Review & Editing, Visualization, Supervision. Elias N. Pergantis: Methodology, Formal Analysis, Investigation, Writing - Original Draft, Writing - Review & Editing. Levi D. Reyes Premer: Investigation, Writing - Review & Editing. Panagiotis Papageorgiou: Investigation, Writing - Review & Editing. Alex H. Lee: Investigation, Writing - Review & Editing. James E. Braun: Writing - Review & Editing. Gregor P. Henze: Conceptualization, Writing - Review & Editing. Kevin J. Kircher: Conceptualization, Methodology, Validation, Writing - Review & Editing, Supervision, Project Administration, Funding Acquisition.
@@
Data are publicly available in [20].